# Anosov diffeomorphisms of open surfaces

### v1

@@ Abstract.
We study the existence of Anosov diffeomorphisms on complete open surfaces. We show that under the assumptions of density of periodic points and uniform geometry that such diffeomorphisms have a system of Margulis measures, which are a holonomy invariant and dynamically invariant system of measures along the stable and unstable leaves. This shows that there can be no such diffeomorphism with a global product structure.
@@ 1. Introduction
A diffeomorphism $f\colon M\to M$ of a closed Riemannian manifold is called Anosov if it preserves a continuous invariant splitting $TM=E^{u}\oplus E^{s}$ and $Df$ uniformly expands vectors in $E^{u}$ and uniformly contracts vectors in $E^{s}$. These bundles are called the stable and unstable bundles of $f$ and are independent of the particular choice of metric on $M$. Anosov diffeomorphisms are of perennial interest in smooth dynamics as they are one of the most fundamental examples of a dynamical system with chaotic behavior. A long standing question of Smale asks if every Anosov diffeomorphism of a closed manifold is a diffeomorphism of a manifold finitely covered by a manifold homeomorphic to a compact quotient of a nilpotent Lie group [Sma67]. A resolution of this question would clarify the relationship between hyperbolic behavior and the geometry and topology of the underlying manifold.
In this paper, we consider the existence of Anosov diffeomorphisms on non-compact Riemannian surfaces. We shall make this notion precise, but informally we mean a diffeomorphism $f\colon M\to M$ of a complete Riemannian surface, such that $f$ preserves an invariant hyperbolic splitting. We study such maps under two assumptions. The first assumption asserts that the periodic points of $f$ are dense, which implies that the stable and unstable manifolds of periodic points are dense. Our second assumption is that the geometry of $M$ and $f$ is uniform in the sense explained by Definition 2.2. Our main result is that such a diffeomorphism has a system of “Margulis measures” defined on its stable and unstable leaves.
@@
One reason it is interesting to construct Margulis measures is that they have strong rigidity applications. For instance, Hiraide gave an elegant proof that a codimension 1 Anosov diffeomorphism is a diffeomorphism of a torus using Margulis measures [Hir01]. See [BFT22] for a more recent construction and application to Anosov flows. Here we illustrate their use with the following application.
Theorem 1.2.
Suppose that $f\colon M\to M$ is a $C^{1}$ uniform Anosov diffeomorphism of a surface such that periodic points are dense and such that the unstable and stable foliations of $M$ have global product structure in the universal cover of $M$. Then $M$ is closed.
The global product structure referred to in the above theorem means that in the universal cover of $M$ each leaf of the lift of the stable foliation meets each leaf of the lift of the unstable foliation in exactly one point.
The proof of Theorem 1.2 appears in Section 5. See [Ham14] for a more detailed discussion of the role of product structure in the classification of Anosov diffeomorphisms.
@@ 1.1. Motivation
While our results give only a partial answer this question, developing tools to address this question is one of the main motivations for our work.
1.2. Examples of Anosov Diffeomorphisms of Open Surfaces
@@
Let us first consider a simple construction that illumiates the role the metric plays. Consider a translation $T\colon\mathbb{R}\to\mathbb{R}$. As $T$ has no recurrence, it is straightforward to find a metric $g$ on $\mathbb{R}$ such that with respect to this metric $\|DT\|>c>1$: for example, take $g=e^{-t}dx^{2}$. However, for any metric on $\mathbb{R}$ for which $DT$ is uniformly bounded below, $\mathbb{R}$ cannot be complete as $T$ would then have a fixed point. Of course, this map isn’t “Anosov” in the sense that it has a splitting into two bundles, but it shows that there are issues that arise when one tries to make a map uniformly hyperbolic by changing the metric.
Let us now consider examples on surfaces. There are several constructions of Anosov diffeomorphisms on $\mathbb{R}^{2}$. The most obvious is a hyperbolic linear map. However, there are some substantially more interesting examples. In 1973, White [Whi73] showed that there is a complete metric and an invariant splitting of $T\mathbb{R}^{2}$, which makes the translation $(x,y)\mapsto(x+1,y)$ into an Anosov diffeomorphism. In White’s example, the unstable foliation is comprised of a translation invariant collection of Reeb components. In addition, the metric that White constructs has unbounded curvature. In 1977, Mendes [Men77] studied Anosov diffeomorphisms of $\mathbb{R}^{2}$ and conjectured that any Anosov diffeomorphism of $\mathbb{R}^{2}$ is topologically conjugate either to a translation or a hyperbolic linear map. For about forty years no new work on the conjecture appeared. However, recently it was shown that the Mendes conjecture was not correct. Matsumoto showed that there exists an Anosov diffeomorphism of $\mathbb{R}^{2}$ that is not conjugate to either of these examples [Mat21]. However, work of Groisman and Nitecki shows that the conjecture is true if one assumes that $f$ is the time one map of a flow [GN21]. An important aspect of this work is that it does not use any assumptions on the uniformity of the geometry of the diffeomorphisms being studied.
@@ 1.3. Outline of the Proof
The underlying geometric idea of our proof is based on work of Hiraide [Hir01], who used a similar approach to prove that a co-dimension one Anosov diffeomorphism of a closed manifold is necessarily a diffeomorphism of a torus. Hiraide’s approach is based on studying the conditional measures of the measure of maximal entropy along a one dimensional foliation. In our case, as in Hiraide’s, once one has constructed the conditionals of the measure of maximal entropy, the result follows by geometric considerations as these conditional measures give strong geometric constraints on the manifold. The main technical difficulty in our work is constructing a family of measures on the leaves of the unstable foliation of $f$ that has the same properties as the conditional measure of MME has in the compact case. We must do this knowing that an MME need not even exist in our setting. We refer to these measures we construct as Margulis measures below in order to emphasize one of their most important properties: the dynamics scales them conformally. We will refer to these measures as $\mu^{u}$ to emphasize that they live on leaves of the unstable foliation.
@@
In subsection 4.4, we show how to push the measures just constructed on $\widehat{\Sigma}$ to $M$ without sacrificing any of their properties. The most important property that must be retained is local holonomy invariance. While this property is immediate on the symbolic system, obtaining it for local stable holonomies on $M$—which need not respect the elements of the Markov partition—is more involved. Consequently, it becomes imperative that the harmonic function has been constructed correctly to ensure these properties.
@@
1.4. Comments on the Proof
The main contribution of this paper is our construction of an analog of the conditional of the Margulis measure in the non-compact setting: we construct a family of measures defined on each unstable leaf such that these measures are invariant on the nose under local stable holonomy in addition to a number of other natural properties, such as being fully supported. Many of these features are unique to the Margulis measure even in the compact case. The construction we pursue uses a coding of the dynamics on $M$ as a countable Markov shift. This approach is not straightforward because there is no longer a canonical Margulis measure on the countable Markov shift that has all the properties we need as holds in the compact case. In fact there are many such measures, and they might even be atomic. Consequently, one need not only carefully construct the correct Margulis measure on the shift-space, but must then push this measure forward onto $M$. As the coding of the system is not unique and the measures we construct potentially have atoms, pushing forward the measure and getting something with usable properties is non-trivial.
One novelty of this work and its approach is that it is a rigidity result that not only avails itself of the recent work on thermodynamic formalism of countable Markov shifts but crucially makes use of this work. It is difficult to see how one could approach the problem without encountering many of the difficulties found here. These difficulties manifest when one passes to the symbolic system and are tied to specific and it seems unavoidable questions about the recurrence or transience of a dynamical system on a non-compact manifold. Contending with such issues concerning recurrence is a fundamental feature of dynamics in a non-compact setting and a main technical difficulty in this work. Additionally, this work seems to be one of the first to make use of countable Markov shifts for studying dynamics in an explicitly non-compact setting.
@@ 1.4. Comments on the Proof
2. Bounded and Uniform Geometry
We will state the question for an Anosov diffeomorphism of an open surface. In order to do this we will formulate a notion of an Anosov diffeomorphism on an open Riemannian surface with geometric qualities that permit analysis without unnecessary technical obstacles.
Definition 2.1.
If $f\colon M\to M$ is a diffeomorphism, then we say that $f$ is Anosov if there exists a continuous metric on $M$ and a continuous $df$-invariant splitting $TM=E^{u}\oplus E^{s}$ such that the following holds. There exist $C,\lambda>0$ such that for all $v\in E^{u}\setminus\{0\}$ and all $w\in E^{s}\setminus\{0\}$,
for all $n\in\mathbb{N}$.
In the case that $M$ is compact, it turns out that being an Anosov diffeomorphism is independent of the metric and can be characterized without reference to any metric. In the non-compact case there is more subtlety .
Definition 2.2.
  1. (1)
  2. (2)
    The angles between the subspaces in the Anosov subspaces are uniformly bounded.
  3. (3)
  4. (4)
  5. (5)
    The surface $M$ has bounded geometry whenever $M$ is smooth enough for this to be defined. If the regularity of $M$ is too low for $M$ to have bounded geometry, then we require that $M$ is a uniform manifold of order one, as described below.
Remark 2.3.
As in the compact setting, under the assumptions of Definition 2.2, one can deduce that uniform hyperbolicity implies that the angle in (2) is uniformly large. However, we leave this as an assumption in the definition for clarity.
The notion of what it means for $M$ to have bounded geometry deserves some discussion. The usual definition of bounded geometry essentially requires two things: positive injectivity radius and bounds on the covariant derivative of the curvature tensor. This definition is fine, but restricts one to studying manifolds that are at least $C^{2}$ with a $C^{2}$ metric. When studying dynamics, this often represents an artificial restriction on the results one can prove. In the case of low regularity, one can instead work with uniform manifolds, we now quote a definition from Eldering [Eld17]:
Definition 2.4.
  1. (1)
    $B(\phi_{i}(x),\delta)\subset\phi_{i}(U_{i});$
  2. (2)
In [Eld17], Eldering shows that, in a sense, manifolds with bounded geometry and uniform manifolds are the same thing.
@@ 3. Markov Partition for Non-Compact Anosov Diffeomorphisms
The purpose of this section is to construct a Markov partition for the dynamics of $f$ on the non-compact space $M$, following Bowen’s construction as it appears in [BR75]. The construction is general, and only requires the properties of bounded geometry. It does not rely on the fact that $\mathrm{dim}M=2$.
@@ 3.1. Basic Properties of Uniform Anosov Diffeomorphisms
    $\|d_{x}f^{n}e^{u}\|\geq Ce^{-\lambda n}\text{ and }\|d_{x}f^{-n}e^{s}\|\leq Ce^{-\lambda n}.$
@@
    Let $r_{0}$ denote the supremum of distance to the boundary of all uniform charts.
@@
(Smale Bracket) Suppose that $f$ is a uniform Anosov diffeomorphism. For sufficiently small $\epsilon>0$, there exists $\delta>0$ such for any $x,y$ with $d(x,y)<\delta$ $W^{u}_{\epsilon}(x)\cap W^{s}_{\epsilon}y)$ consists of a single point denoted $[x,y]$. Further the map $[\cdot,\cdot]\colon(x,y)\mapsto[x,y]$ defined on pairs of points $(x,y)$ where $d(x,y)<\delta$ is uniformly continuous. This map is called the Smale bracket.
@@
In our main results, we assume that the periodic points of $f$ are dense in $M$. Due to the existence of local product neighborhoods for uniform Anosov diffeomorphisms, it is straightforward to show that this implies that the stable and unstable manifolds of each periodic point are dense in $M$. Consider the closure $C$ of an (un)stable leaf. Any open ball near the boundary of $C$ contains a periodic point whose stable manifold intersects $C$. This leads to a contradiction.
@@ 3.2. Background on Countable Markov Shifts
Let $S$ be a countable collection of states. And let $(\mathbf{A}_{ij})_{S\times S}$ be a (infinite) $\{0,1\}$-matrix. We define the countable Markov shift as the dynamical system on the set
$\widehat{\Sigma}=\{\omega\in S^{\mathbb{Z}}:A_{\omega_{n}\omega_{n+1}}=1\text{for all }n\in\mathbb{Z}\}.$
@@
For a single symbol $R\in S$, we define the cylinder set
@@ 3.3. Induced topological Markov shift from a Markov Partition
We begin by introducing the shift space and its associated coding map. The proof of the following is the same as [Bow08, Theorem 3.18] and [Bow08, Lem. 4.2].
@@
That this definition makes sense and has the needed properties is the content of the following theorem.
Theorem 3.14.
Proof.
That the map is well-defined and Hölder continuous is the same as in the compact case, so we omit this. We check that $\widehat{\pi}$ is uniformly finite to one below in Proposition 3.20.
We just check surjectivity. Let $x\in M$; we show $x\in\widehat{\Sigma}$. As $Y$ is dense, there is a sequence $x_{n}\in Y$ such that $x_{n}$ converges to $x$. Let $\underline{R}^{(n)}$ be the preimage of $x_{n}$ under $\hat{\pi}$. Then as the Markov partition is locally finite, we may assume that $R_{0}^{(n)}=R$ for some fixed $R\in\mathcal{R}$. Thus by local compactness of $\widehat{\Sigma}$, we may assume that $\underline{R}^{(n)}\to\underline{R}$ for some $\underline{R}\in\widehat{\Sigma}$. Thus by continuity of $\hat{\pi}$, $\hat{\pi}(\underline{R})=x$. ∎
Later we will make a close study of the points in $M$ such that $\left|\widehat{\pi}^{-1}(\{x\})\right|=1$. Of particular importance is the following set.
@@
Let $Y^{\prime}$ denote the set of points $x$ whose stable leaves never meet a stable boundary of a rectangle and whose unstable leaves never meet an unstable boundary of a rectangle. Let $Y^{\prime}$ denote this other set. Note that the set $Y^{\prime}$ is closed under the Smale bracket.
@@
Remark: Inside each standard unstable leaf $W^{u}(x,R)$, the meagre set which is complement to $Y$ is countable. The number of unstable leaves which are not standard is countable.

### v2

@@ Abstract.
We study the existence of Anosov diffeomorphisms on complete open surfaces. We show that under the assumptions of density of periodic points and uniform geometry that such diffeomorphisms have a system of Margulis measures, which are a holonomy invariant and dynamically invariant system of measures along the stable and unstable leaves.
@@ 1. Introduction
A diffeomorphism $f\colon M\to M$ of a closed Riemannian manifold is called Anosov if it preserves a continuous invariant splitting $TM=E^{u}\oplus E^{s}$ and $Df$ uniformly expands vectors in $E^{u}$ and uniformly contracts vectors in $E^{s}$. These bundles are called the stable and unstable bundles of $f$ and are independent of the particular choice of metric on $M$. Anosov diffeomorphisms are of perennial interest in smooth dynamics as they are one of the most fundamental examples of a dynamical system with chaotic behavior.
In this paper, we consider the existence of Anosov diffeomorphisms on non-compact Riemannian surfaces. We shall make this notion precise, but informally we mean a diffeomorphism $f\colon M\to M$ of a complete Riemannian surface, such that $f$ preserves an invariant hyperbolic splitting. We study such maps under two assumptions. The first assumption asserts that the periodic points of $f$ are dense, which implies that the stable and unstable manifolds of periodic points are dense. Our second assumption is that the geometry of $M$ and $f$ is uniform in the sense explained by Definition 1.3. Before we proceed we will make precise the setting that we are considering.
1.1. Definition of Uniform Anosov Diffeomorphisms
We now formulate a notion of an Anosov diffeomorphism on an open Riemannian surface with geometric qualities that permits analysis without unnecessary technical obstacles.
Definition 1.1.
If $f\colon M\to M$ is a diffeomorphism of a Riemannian manifold $M$, then we say that $f$ is Anosov if there is a continuous $df$-invariant splitting $TM=E^{u}\oplus E^{s}$ such that the following holds. There exist $C,\lambda>0$ such that for all $v\in E^{u}\setminus\{0\}$ and all $w\in E^{s}\setminus\{0\}$,
for all $n\in\mathbb{N}$.
In the case that $M$ is compact, it turns out that being an Anosov diffeomorphism is independent of the metric and can be characterized without reference to any metric. In the non-compact case there is more subtlety. For this we have the notion of a uniform manifold, which is a convenient generalization of the notion of bounded geometry.
[…]
Definition 1.2.
  1. (1)
    $B(\phi_{i}(x),\delta)\subset\phi_{i}(U_{i});$
  2. (2)
We call these charts uniform charts.
With this notion, we can formulate our setting.
Definition 1.3.
  1. (1)
  2. (2)
    The angles between the subspaces in the Anosov splitting are uniformly bounded below.
  3. (3)
  4. (4)
  5. (5)
    $M$ is a uniform manifold.
Note that a manifold supporting a uniform Anosov diffeomorphism comes with an atlas of uniform charts.
Remark 1.4.
As in the compact setting, one can deduce (2) Definition 1.3 from (1), (3), (4), (5). However, we leave this as an assumption in the definition for clarity.
1.2. Main Result
Our main result is that a uniform Anosov diffeomorphism has a system of “Margulis measures” defined on its stable and unstable leaves as long as it has sufficiently many periodic points.
@@ 1.3. Motivation
While our results give only a partial answer this question, developing tools to address this question is one of the main motivations for our work as the construction of Margulis measures can help address this problem.
Hiraide gave an elegant proof that a codimension 1 Anosov diffeomorphism is a diffeomorphism of a torus using Margulis measures [Hir01]. See [BFT22] for a more recent construction and application to Anosov flows. Here we illustrate their use with the following application.
Theorem 1.6.
Suppose that $f\colon M\to M$ is a $C^{1}$ uniform Anosov diffeomorphism of a surface that admits a system of Margulis measures as in Theorem 1.5 giving infinite mass to each unstable leaf. Then $M$ is closed.
In particular, if one could construct such Margulis measures for invariant leaves of the lamination appearing in 1, it would give an affirmative answer to the question.
1.4. Examples of Anosov Diffeomorphisms of Open Surfaces and Manifolds
@@
Let us first consider a simple construction that illuminates the role the metric plays. Consider a translation $T\colon\mathbb{R}\to\mathbb{R}$. As $T$ has no recurrence, it is straightforward to find a metric $g$ on $\mathbb{R}$ such that with respect to this metric $\|DT\|>c>1$: for example, take $g=e^{-t}dx^{2}$. However, for any metric on $\mathbb{R}$ for which $DT$ is uniformly bounded below, $\mathbb{R}$ cannot be complete as $T$ would then have a fixed point. Of course, this map isn’t “Anosov” in the sense that it has a splitting into two bundles, but it shows that there are issues that arise when one tries to make a map uniformly hyperbolic by changing the metric.
Let us now consider examples on surfaces. There are several constructions of Anosov diffeomorphisms on $\mathbb{R}^{2}$ in the sense of Definition 1.1. These examples differ in terms of what types of metrics are put on $\mathbb{R}^{2}$. The most obvious is a hyperbolic linear map, for which the metric has bounded geometry. However, there are some substantially more interesting examples. In 1973, White [Whi73] showed that there is a complete metric and an invariant splitting of $T\mathbb{R}^{2}$, which makes the translation $(x,y)\mapsto(x+1,y)$ into an Anosov diffeomorphism. In White’s example, the unstable foliation is comprised of a translation invariant collection of Reeb components. In addition, the metric that White constructs has unbounded curvature. In 1977, Mendes [Men77] studied Anosov diffeomorphisms of $\mathbb{R}^{2}$ and conjectured that any Anosov diffeomorphism of $\mathbb{R}^{2}$ with respect to a complete metric is topologically conjugate either to a translation or a hyperbolic linear map. For about forty years no new work on the conjecture appeared. However, recently it was shown that the Mendes conjecture was not correct. Matsumoto showed that there exists an Anosov diffeomorphism of $\mathbb{R}^{2}$ with respect to a complete metric that is not conjugate to either of these examples [Mat21]. However, work of Groisman and Nitecki shows that the conjecture is true if one assumes that $f$ is the time one map of a flow [GN21]. An important aspect of this work is that it does not use any assumptions on the uniformity of the geometry of the diffeomorphisms being studied.
@@
Besides surfaces, there is interest in Anosov diffeomorphisms and flows on other manifolds motivated both by Question 1 and understanding Anosov dynamics more generally. There are not many papers that explicitly consider this question from the point of view of classification. In [BL25], a characterization is obtained showing when the lift of $\mathbb{R}$-covered Anosov flow is transitive [BL25, Thm. C]. In particular, this allows them to construct examples of transitive Anosov flows on open manifolds that are lifts of flows on closed manifolds [BL25, Thm. D]. In addition, certain kinds of quantitative results have been proved, for example, for counting geodesics in infinite covers under the assumption of transitivity of the dynamics. See [Pol17] or [DS21]. These results are also related, for example, to certainly physically motivated models. For example, in the study of the $\mathbb{Z}^{d}$-periodic Lorenz gas, the periodicity of the gas is analogous to the dynamics of a hyperbolic flow on a $\mathbb{Z}^{d}$-cover, see for example [DNP22]. In view of the particular applications mentioned above, it is quite natural to consider the existence of transitive examples, as [BL25] does.
The authors are not sure if the techniques here suffice to prove the analogous result for Anosov flows on non-compact $3$-manifolds. A typical approach to constructing Markov partitions of flows makes use of the first return map to a section. For this return map it seems likely that the non-compactness can introduce additional complications beyond those that exist for just diffeomorphisms. We note, however, that for some non-compact manifolds, Margulis measures may be obtained without effort. If we start with an Anosov flow $f^{t}$ on a compact manifold $M$, and we lift $f^{t}$ to an infinite cover of $M$, then the lift will have Margulis measures obtained by lifting those in the base.
@@ 1.5. Outline of the Proof
In subsection 3.5, we show how to push the measures just constructed on $\widehat{\Sigma}$ to $M$ without sacrificing any of their properties. The most important property that must be retained is local holonomy invariance. While this property is immediate on the symbolic system, obtaining it for local stable holonomies on $M$—which need not respect the elements of the Markov partition—is more involved. Consequently, it becomes imperative that the harmonic function has been constructed correctly to ensure these properties. This challenge is similar to the challenge which Bowen and Marcus face in [BM77] when constructing an invariant measure to the horocyclic flow of an Axiom A diffeomorphism on a compact manifold.
@@ 2. Markov Partition for Non-Compact Anosov Diffeomorphisms
The purpose of this section is to construct a Markov partition for the dynamics of $f$ on the non-compact space $M$, following Bowen’s construction as it appears in [BR75]. The construction is general, and only requires uniformity properties. It does not rely on the fact that $\mathrm{dim}\,M=2$.
@@ 2.1. Basic Properties of Uniform Anosov Diffeomorphisms
    $\|d_{x}f^{n}e^{u}\|\geq Ce^{\lambda n}\text{ and }\|d_{x}f^{-n}e^{s}\|\geq Ce^{\lambda n}.$
@@
    Let $r_{0}$ denote the supremum of distance to the boundary of the domain of all uniform charts.
@@
(Smale Bracket) Suppose that $f$ is a uniform Anosov diffeomorphism. For sufficiently small $\epsilon>0$, there exists $\delta>0$ such for any $x,y$ with $d(x,y)<\delta$, $W^{u}_{\epsilon}(x)\cap W^{s}_{\epsilon}y)$ consists of a single point denoted $[x,y]$. Further the map $[\cdot,\cdot]\colon(x,y)\mapsto[x,y]$ defined on pairs of points $(x,y)$ where $d(x,y)<\delta$ is uniformly continuous. This map is called the Smale bracket.
@@
In our main results, we assume that the periodic points of $f$ are dense in $M$. Due to the existence of local product neighborhoods for uniform Anosov diffeomorphisms, it is straightforward to show that this implies that the stable and unstable manifolds of each periodic point are dense in $M$. To see this, consider the closure $C$ of an (un)stable leaf: Any open ball near the boundary of $C$ contains a periodic point whose stable manifold intersects $C$. This leads to a contradiction.
@@ 2.2. Background on Countable Markov Shifts
Let $\mathcal{R}$ be a countable collection of symbols. And let $(\mathbf{A}_{ij})_{\mathcal{R}\times\mathcal{R}}$ be a (infinite) $\{0,1\}$-matrix. We define the countable Markov shift as the dynamical system on the set
$\widehat{\Sigma}=\{\omega\in\mathcal{R}^{\mathbb{Z}}:A_{\omega_{n}\omega_{n+1}}=1\text{ for all }n\in\mathbb{Z}\}.$
For an element $\underline{R}\in\hat{\Sigma}$ we write $R_{i}$ for its $i$th entry, $i\in\mathbb{Z}$.
@@
For a single symbol $R\in\mathcal{R}$, we define the cylinder set
@@ 2.3. Induced topological Markov shift from a Markov Partition
We begin by introducing the shift space and its associated coding map.
@@
The proof that this definition and hence the coding associated to a Markov partition makes sense is the same as the proofs of [Bow08, Theorem 3.18] and [Bow08, Lemma 4.2].
Later we will make a close study of the points in $M$ such that $\left|\widehat{\pi}^{-1}(\{x\})\right|=1$. Of particular importance is the following two sets, $Y$ and $Y^{\prime}$, which we will reference throughout the rest of the paper.
@@
Let $Y^{\prime}$ denote the set of points $x$ whose stable leaves never meet a stable boundary of a rectangle and whose unstable leaves never meet an unstable boundary of a rectangle. Note that the set $Y^{\prime}$ is closed under the Smale bracket.
One of the main technical issues we face is dealing with leaves that lie in the boundaries of rectangles. For this reason we will focus on the leaves that avoid the orbits of the boundaries as much as possible.
@@
Remark: Inside each standard unstable leaf $W^{u}(x,R)$, the meagre set which is the complement to $Y$ is countable. The number of unstable leaves which are not standard is countable.
@@
The following theorem summarizes the main properties of the coding in Definition 2.13. We say that a map $\phi$ is uniformly locally Hölder continuous when there is are fixed $C,\alpha>0$ such that the restriction of $f$ to sets of uniform diameter $\delta>0$ is $\alpha$-Hölder with Hölder constant $C$.
Theorem 2.20.
Proof.
That the map is well-defined and Hölder continuous is the same as in the compact case, so we omit this. That $\widehat{\pi}$ is uniformly finite to one was checked in Proposition 2.19.
We just check surjectivity. Let $x\in M$; we show $x\in\widehat{\Sigma}$. Let $Y$ be as in Definition 2.14. As $Y$ is dense, there is a sequence $x_{n}\in Y$ such that $x_{n}$ converges to $x$. Let $\underline{R}^{(n)}$ be the preimage of $x_{n}$ under $\hat{\pi}$. Then as the Markov partition is locally finite, we may assume that $R_{0}^{(n)}=R$ for some fixed $R\in\mathcal{R}$. Thus by local compactness of $\widehat{\Sigma}$, we may assume that $\underline{R}^{(n)}\to\underline{R}$ for some $\underline{R}\in\widehat{\Sigma}$. Thus by continuity of $\hat{\pi}$, $\hat{\pi}(\underline{R})=x$. ∎